8b.A projectile is launched with an initial speed of 59.0 m/s at an angle of 35.0° above the horizontal. The projectile lands on a hillside 3.55 s later. Neglect air friction. (Assume that the +x-axis is to the right and the +y-axis is up along the page.)

(a) What is the projectile's velocity at the highest point of its trajectory?


(b) What is the straight-line distance from where the projectile was launched to where it hits its target?

 


1
Expert's answer
2022-06-06T08:54:23-0400

(a)

"v_x=v_{0x}=v_0\cos\theta\\
=59.0*\cos35.0^\circ=48.3\:\rm m/s"

"v_y=0"

(b)

"x=v_0\cos\theta t=\\
=59.0*\cos35.0^\circ*3.55=172\:\rm m"

"y=v_0\sin\theta t-gt^2/2=\\
=59.0*\sin35.0^\circ*3.55-9.8*3.55^2/2=58.4\:\rm m"

"d=\sqrt{x^2+y^2}=\sqrt{172^2+58.4^2}=182\:\rm m"